OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (anonymous):

15x-30y=15 42x-30y=-66 Now, -30y=-66+42x So 15x-66+42x =15 57x-66=15 57x=15+66=81 57x=81 x=81/57 Now, 15(81/57)-30y=15, solve for y

OpenStudy (anonymous):

It's probably easier though if you make the x values equal

OpenStudy (anonymous):

so 21x-42y=21 21x-15y=-33

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

im really not good at eliminating

OpenStudy (anonymous):

21x-42y=21 21x-15y=-33 Solve for y: 21x = -33+15y So, -33+15y-42y=21 Solve for y

OpenStudy (anonymous):

would i add 33 to both sides?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sorry, 27y =54

OpenStudy (anonymous):

divide both by 27

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is two but the only answer for it is -2

OpenStudy (anonymous):

its -27y=54

OpenStudy (anonymous):

15y-42y=-27y

OpenStudy (anonymous):

-33+15y-42y=21 Solve for y becomes: -27y = 54 y=-2

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome, good luck